OpenStudy (anonymous):

If you throw a dice 6 times, what's the chance that you'd get a six on: a: exactly one of the throws. b: one or more of the throws.

OpenStudy (amistre64):

thats a pretty big sample space :)

OpenStudy (anonymous):

? I don't understand what your saying... lol

OpenStudy (amistre64):

1 6 15 20 15 6 1 neither do i yet :)

OpenStudy (anonymous):

a. (1/6)*(5/6)^5*6 = (5/6)^5 b. (5/6)^6

OpenStudy (anonymous):

LOL! I was a fool in my math class so I had extra homework... ( threw an apple at the teacher)

OpenStudy (anonymous):

er, b. 1-(5/6)^6

OpenStudy (amistre64):

P(6 a / a / a / a / a / ) = P(6)*P(\)^5, where / means not a 6

OpenStudy (amistre64):

P(6) = 1/6; P(/) = 5/6

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

satellite is good at these, he does them instead of the crosswords during breakfast :)

OpenStudy (anonymous):

exactly one: \[\dbinom{6}{1}\frac{1}{6}\times (\frac{5}{6})^5\]

OpenStudy (amistre64):

he has rooms filled with unthrown dice just waiting to be explored lol

OpenStudy (anonymous):

they are thrown.

OpenStudy (anonymous):

since 6 choose 1 is 6, this answer is really \[(\frac{5}{6})^5\]

OpenStudy (anonymous):

Dang! I didn't think I would get so much attention :P

OpenStudy (anonymous):

one or more throws means not no sixes. the probability you get no sixes is \[(\frac{5}{6})^6\] so your answer is \[1-(\frac{5}{6})^6\]